LOG-IN
Displaying reports 1-25 of 36311.Go to page 1 2 3 4 5 6 7 8 9 10 End
AdV-DET (Commissioning)
Print this report.
genin, fiori - 11:01, Saturday 23 September 2017 (39564)Get code to link to this report
Comment to Investigations on OMC angular noise coupling and best SDB1 angular set points (Click here to view original report: 39537)

Regarding the 270 Hz peak which seems to be present on DARM since the Laser room to SIB2 room door and cleanroom to SIB2 door have been opened. We can clearly see it on the 

the Beam Pointing control system signals (see attached plot). We could not see any excess of accoustic noise at that frequency but there is a slight effect on the seismic noise measured with the accelerometer on top of the EIB which suggest that a resonance of a mount (maybe EIB_M6 one) could be excited (see entry https://logbook.virgo-gw.eu/virgo/?r=38640).

The strange thing is that the excess of noise read by the BPC quadrants is very small and we are not expecting any limitation from input beam jitter noise at that level when looking at the last noise projection we made on the 13th of september (https://logbook.virgo-gw.eu/virgo/?r=39459). At that time, we injected a large boradband noise without being able to see any coherence for frequencies higher than 100Hz.

Maybe this noise comes from the CB clean rooms and couple the ITF through PRCL (SIB2 PDs)?

We plan to make more investigations next week on that issue trying to understand it better and solve it.

Images attached to this comment
39564_20170923110048_plot3.gif
AdV-COM (AdV commissioning (1st part) )
Print this report.
ruggi - 00:03, Saturday 23 September 2017 (39561)Get code to link to this report
Tuning of alignment set poit and gains

The alignment has been improved changing the set point of some quadrants. The tuning has been effective for coming back close to 30 Mpc. The 270 Hz peak is still there but slightly lower.

A good control filter has been found for DIFFp_TX. This works fine in the present quiet conditions, but probably is not the best one when the seismic noise is high. Anyway, the accuracy we got tonight is the best ever achieved.

Images attached to this report
39561_20170923000441_atun1.jpg 39561_20170923000449_atun3.jpg 39561_20170923000458_atun2.jpg 39561_20170923000504_atun.jpg
Detector Operation (Operations Report)
Print this report.
Gherardini - 23:01, Friday 22 September 2017 (39559)Get code to link to this report
Operator Report - afternoon shift 22/09
This shift was dedicated to alignment, the activity went on without any major problem, activity in progress...

-automation
because of some mistake/typo we had to change a parameter in the INJ_MAIN.ini file, using the "LOAD" button on the client was not enough I had to stop & start the INJ_MAIN metatron node to make the modification effective.

-inf
trying to rise up the laser lab temperature that after the problem of the air conditioning system is decreased of 1 degree at around 12:00UTC I closed half door between the laser lab and the SIB2 room, up to now the action seems to not have a big effect (see plot).

Images attached to this report
39559_20170922230110_0922injac.gif
Mirror Suspension Control (Mirror Suspension Control)
Print this report.
ruggi - 19:54, Friday 22 September 2017 (39553)Get code to link to this report
Effect of F7 TX and TZ dampers on DARM

We found that the TX and TZ dampers of suspension filter 7 are injecting control (sensing) noise  on DARM between 2 and 6 Hz. This is evident only since DARM has been improved in that region due to the mitigation of PR_MAR_TX control noise. The largest contribution comes from WI. The marker of that is the presence of two narrow peaks around 3 Hz in WI_F7 sensing, which are well visible in DARM and disappear (together with some broadband noise) when the dampers are turned off. DARM improves more when the dampers of all the suspensions are turned off (Fig. 1). A precise noise budget of each d.o.f. has not been done.

We tested a lock acquisition with all the dampers switched off at the beginning, but some oscillation started to grow up when the dark fringe was reached. We moved the disabling of the dampers just before the end of the locking procedure, and we saw no bad effec for a few hours. But later in the night the NE payload got excited, mainly at 0.4 Hz along the transversal axis, and the effect was visible in the sensitivity at low frequency. It would be interesting to make some specific test in order to understand the coupling mechanism, and if it is present on all the suspension or only on NE.

Tonight I'm going to leave the ITF with all the dampers disabled except NE, in order to see if some other slow growing of resonances shows up.

Images attached to this report
39553_20170922192751_f7damp.jpg
AdV-DET (Commissioning)
Print this report.
bersanetti, ruggi, gouaty, bonnand, frolov, kasprzack - 19:43, Friday 22 September 2017 (39554)Get code to link to this report
Comment to Investigations on OMC angular noise coupling and best SDB1 angular set points (Click here to view original report: 39537)
We wanted to investigate if the 270 Hz and nearby noise peaks are from the OMC beam jitter.
The first plot compares the best reference from last week (when the 270 peaks were very low) to the current typical noise floor and to the two bench misalignments (TX, TY). In one of the misaligned state (TY), the 270 peak height increased as if it was driven by the beam jitter. But we didn't observe the same increase for the other misalignments - TX or TY.
The second plot shows the response to excitations from 100 to 400 Hz. A different set of peaks was excited while the height of 270 peaks remains unaffected. This would mean that we are not only looking at OMC beam jitter coupling, but also at some other noise coupling that gets affected by the global alignment (by moving SDB1, we are also moving the beam on SDB2, which has QPDs used for common arm alignment).

Furthermore, the DARM coherence with PRCL at 270 Hz is high according to BRUCO (see third plot). This makes us suspect that the peak is actually from frequency noise coupling, and it gets modulated by the global alignment.

The 270 peak was persistently in the DARM spectrum starting from September 15. Prior to this date, there were several days when the best spectrum of the day didn't have the peak. Coincidentally, the injection lab AC status changed on September 14.

During the same shift, we tried to see if SPRB alignment had an effect on the peaks, but scanning the full SPRB_B4_112 and LSC_B4_DC ranges didn't modulated the 270 peak.
Images attached to this comment
39554_20170922193531_sdb1tyoffset.png 39554_20170922193539_sdb1exc.png 39554_20170922193548_lscprcl.png
AdV-INF (INJ Lab re-arrangements)
Print this report.
derossi, genin - 18:22, Friday 22 September 2017 (39552)Get code to link to this report
Comment to Failure of the INJ area ACS and workaround (Click here to view original report: 39541)

The decreasing of the IMC transmitted power is probably a consequence of  the decreasing of the pre-mode cleaner transmitted power, due to a misalignment of the cavity linked to the temperature.

Images attached to this comment
39552_20170922181931_pmclosspower.png
AdV-COM (AdV commissioning (1st part) )
Print this report.
ruggi - 17:55, Friday 22 September 2017 (39551)Get code to link to this report
Comment to Many changes in suspension and alignment control strategy (Click here to view original report: 39493)

The problem of fModErr loop has been fixed yesterday evening. Now GIPC can be used without affecting fModErr.

AdV-COM (AdV commissioning (1st part) )
Print this report.
fiori, pasqualetti, paoletti, chiummo, vacuum - 17:50, Friday 22 September 2017 (39550)Get code to link to this report
mechanical damping of B1 mini-link

Following up the tests performed on Sept 5th (39382), this morning the vacuum team worked on B1 mini-link aiming to dump its mechanical vibrations. Each of the two bellows has been wrapped with one layer of silicon rubber (2mm thick) placed in contact with the pipe and on top of it one layer of lead (1mm thick sheet) the whole gently tightened with a metal strip, see pictures 1 (before) and 2 (after).

With one mini-accelerometer sticked to the pipe (sensing in the direction perpendicular to the pipe axis) we measured the mechanical effect of this installation, the plots refer to quiet noise condition:

  • before (blue line in Figure 3) the first mode of the link was around 51Hz, after it moved a bit down at around 44Hz and it reduced in amplitude (Fig. 3 purple line)
  • other peaks around 84Hz, 255Hz (triplet) and 405 (doublet) Hz are no more visible in the quiet vibration spectrum (see also Figure 4 and Fig 5, for zoom in)

We also tried tapping the link before and after the istallation: while before the tapping immediately excited one resonant mode seen by the GEO on SDB2 F0 (Figure 6) in particular a resonance at 52Hz which occasionally happened to match the pipe resonance (as foreseen by A.Bertolini in 39397) eventually saturating the geophone signals, in the new configuration the resulting excitation of the suspension measured by the geophones is less pronounced and did not saturate (Figure 7). However this test was better repeated with the ITF locked, as described in 39547.

Finally, we also measured the resonant modes of the other pipes, for characterization purposes (that is we sticked the mini-acc on each one in turn and measured a quiet spectrum). We annotate the frequency of the first mode:

  • B1p pipe (Figure 8): 30 Hz
  • HS beam pipe (Fig. 9):  24 Hz
  • B5 pipe (Fig. 10): 45 Hz
Images attached to this report
39550_20170922174622_img20170921114824.jpg 39550_20170922174626_img20170922114530.jpg 39550_20170922174634_result.gif 39550_20170922174642_result1.gif 39550_20170922174647_result2.gif 39550_20170922174656_tappinglinkb11.gif 39550_20170922174706_tappinglinkb1config21.gif 39550_20170922174726_quietolinkb1pzoom.gif 39550_20170922174734_quietolinkhsbeam.gif 39550_20170922174740_quietolinkb5.gif
AdV-INF (INJ Lab re-arrangements)
Print this report.
dattilo - 17:33, Friday 22 September 2017 (39549)Get code to link to this report
Comment to Failure of the INJ area ACS and workaround (Click here to view original report: 39541)

In Figs 1 and 2 is shown that starting from the time when the pressure probe broke down (PRES_OUT channel) and the fan speed (FREQ_OUT channel) increased at the maximum value, the INJ system started to have a lot of  unlocks  (IMC_TRA_DC channel).  The origin of the problem was not immediately recognized as an ACS problem, because the lab temperature remained constant (LLR_TE channel).  This caused problems during  the Thursday 14th afternoon shift.  I will suggest to the ACS referent person some improvements in the alarms criteria such that the DMS can recognize and alert on similar kind of plant failures.

Concerning the effect of a different working point of the temperature in the Laser Lab after the workaround, in Fig3 is reported a longer trend of IMC power, were it is possible that it  decreased as consequence of the ACS failure.

Images attached to this comment
39549_20170922173511_injfailure14sept2017imcpower2.gif 39549_20170922173517_injfailure14sept2017imcpower2zoom.gif 39549_20170922173524_imctrapowertrend.gif
Interferometer_sensitivity_studies (General)
Print this report.
Paoletti, Fiori, Pasqualetti - 16:15, Friday 22 September 2017 (39547)Get code to link to this report
Comment to Tapping tests in DET Lab -- part 2 (Click here to view original report: 39382)

Today, after the damping of the N-3 link tube done by Vac team (it will be reported in another dedicated entry), we repeated tapping tests while interferometere was locked.

We did not have the same problems reported in entry 39373 (SDB2 out of control and interferometer unlocked due to SDB2 IP Geophones saturation), so the work done on the link is an improvement.

We report the GPS time (UTC) of the tapping tests, with some brief comments.

- please have a look at linked entry for figures and explanation of links location

(note: we move the little temporary accelerometer ENV_SDB2_LINK_ACC over every “tapped” link)

 

tapping over “N” Link

Time: Sep22,17-12:35:33 -> GPS: 1190118949

Time: Sep22,17-12:36:36 -> GPS: 1190119012

We see a peak around 90Hz in the accelerometer, and (maybe) also in Hrec.

Also, some structures at 240Hz in Hrec

 

tapping over N -1

Time: Sep22,17-12:39:56 -> GPS: 1190119212

Time: Sep22,17-12:40:56 -> GPS: 1190119272

Low frequency (<30Hz) noise in Hrec

 

tapping over N-2

Time: Sep22,17-12:44:09 -> GPS: 1190119465

Time: Sep22,17-12:45:33 -> GPS: 1190119549

We see structures between 60 and 70Hz in Hrec

 

tapping over N-3

Time: Sep22,17-12:48:39 -> GPS: 1190119735

Time: Sep22,17-12:50:21 -> GPS: 1190119837

No visible structures in Hrec, moreover the SDB2 is now stable (interferometer does not unlock as before)

 

accelerometer ENV_SDB2_LINK_ACC is left over this last link (N-3) for the time being

Detector Operation (Operations Report)
Print this report.
Montanari - 15:59, Friday 22 September 2017 (39545)Get code to link to this report
Operator Report - Morning Shift 22/09

This morning I found ITF locked in LOW_NOISE_3 with Horizon ~22MPc.
The Horizon went down in according with the increase in sea activity of this night (see plot). This is a consequence of test of yesterday evening (F7 dampers opened on WI and NI). V.Boschi provide to reclose them at 07:45UTC.

At 08:12UTC I unlocked ITF in according with starting of tapping test around SDB2 and SDB1. In order to allow this activity, I open the SDB2_LC and SDB2_SBE loops.

At 10:10UTC, first part of tapping test, with ITF unlocked, finished, so I closed the SDB2 loops and relocked the ITF (LOW_NOISE_3 at 10:29UTC).
After lunch, work in DET lab continued until 13:00UTC with ITF locked.

After that, activities of afternoon shift (ASC) started.


HVAC
08:14UTC-10:10UTC & 12:14UTC-13:05UTC
In order to allow the work of people inside the DET lab I switched the conditioning system from PR "Portata Ridotta" (Lowered Flow Rate) to PN "portata nominale" (Nominal Flow Rate).

VPM
InfraMoni stopped and restarted at 09:27UTC, under request of I.Fiori I added two flags to monitor the air conditioning flux in DET Lab (HVAC_DET_DLAB_FLUX_IN, HVAC_DET_DLAB_FLUX_OUT).

Images attached to this report
39545_20170922145757_22sepnight.gif
AdV-TCS (TCS studies sensing (simulations + experiments))
Print this report.
rocchi - 11:53, Friday 22 September 2017 (39544)Get code to link to this report
Sidebands stability for different modulation frequencies

In different occasions, Benoit suggested to use a modulation frequency of 18 MHz to improve the SNR of the locking and alignment signals.

One of the objections raised was that the 18 MHz sideband may be as sensitive to aberrations as the 6 MHz.

For this reason, simulations have been performed, scanning the PR RoC and looking at the gain for sidebands at different modulation frequencies.

The results are plotted in Fig. 1 (All modes) and Fig. 2 (TEM00 mode). As it can be seen, indeed the 18 MHz shows a large sensitivity to aberrations and it seems there would be no gain in using such modulation frequency.

Images attached to this report
39544_20170922115144_170922compare61856nomapsnotlall.png 39544_20170922115204_170922compare61856nomapsnotltem00.png
AdV-INF (INJ Lab re-arrangements)
Print this report.
carbognani, dattilo - 11:49, Friday 22 September 2017 (39541)Get code to link to this report
Failure of the INJ area ACS and workaround

As mentioned in this operator report  on Friday 15th  we were forced to switch off the air conditioning system of the INJ area due to a failure.

Here we report a few details on the failure and a check of related side effects.

The sequence of the failure is the following (Fig1 and zoom in Fig2, red plots):

- on Thursday 14th at 14:15 UTC a pressure sensor broke down (channel INJ_PRES_OUT) and its  value dropped to 0;

- this triggered the correction of the pressure loop to increase from ~10% to the maximum value (100%);

- this required the increase of the output  fan speed (channel INJ_FREQ_OUT) from ~22Hz to 60Hz to increase the air flux (channel  INJ_FLUX_OUT) and pressure;

- this stressed the air ducts that started to leak along a few joints;

- the maximum speed was enough to compensate the leakages and keep the system at the nominal temperatures, so the failure was not immediately spotted;

- the day after (Fri 15th) the problem was detected due to the fact that the leakages and the maximum speed were producing disturbances to the operation of the INJ system.

- during the morning  of Fri 15th a temporary repair (also with an external firm) of the most evident leaks was attempted, but not sufficient because other leaks are present further down on the air ducts, some of them in areas not easily accessible. Once  it appeared evident that the  repair is  complicated and time consuming, on the late morning the system was switched off (because it was producing too many IMC unlocks).

- on Friday afternoon a temporary solution was adopted to cope with this issue:  at 15:30 UTC the INJ lab  was put in communication (by opening a door) with the CEB Clean Rooms to take advantage of its air conditioning system that is independent and in operation.  We are still in this configuration that seems enough to avoid downtimes and ensure the scheduled commissioning activities.

Looking at Figs 1 and 2 it's evident that the stress generated on the air ducts by staying at fans maximum speed it's not to be considered anomalous because when in nominal flux condition (channel INJ_MODE_SET value set from 2 to 1) the speed of the fan is comparable. In Fig3 it can be seen that the fan speed, when in nominal state, is at ~50Hz (sept 2016) or ~55Hz (sept2015), not much lower with respect to the maximum one (60Hz), occurred at the time of the failure.

We think that  this accident has simply accelerated a latent problem on the ducts that need to be fully investigated.

In Figs 4 and 5 are reported trends of a few signals were it is possible to see that the temporary solution above described is ensuring acceptable values of temperature fluctuations (although around another working point), humidity, dust contamination (Fig5) and  environmental noise. In some frequency ranges the noise  is a bit higher, in others (example in Figs 6 and 7) it seems even lower.

Concerning the dust contamination we have to check the situation when people are working in the laser lab; however if the workaround solution has to last for several days, we can bring a portable air blower (currently in TCS room) in the laser lab and switch on it  when persons are working there.

Futher checks and inspection of the ducts are planned for the next Tuesday morning maintenance .

Images attached to this report
39541_20170922102806_injfailure14sept2017.gif 39541_20170922102912_injfailure14sept2017zoom.gif 39541_20170922111927_injfailure14sept2017freq.gif 39541_20170922112351_overallinjlab.jpg 39541_20170922112425_injdust.jpg 39541_20170922112454_spectrobsxty010604800.jpg 39541_20170922112504_spectroenvlbseisz060604800.jpg
Comments related to this report
dattilo - 17:33, Friday 22 September 2017 (39549)

In Figs 1 and 2 is shown that starting from the time when the pressure probe broke down (PRES_OUT channel) and the fan speed (FREQ_OUT channel) increased at the maximum value, the INJ system started to have a lot of  unlocks  (IMC_TRA_DC channel).  The origin of the problem was not immediately recognized as an ACS problem, because the lab temperature remained constant (LLR_TE channel).  This caused problems during  the Thursday 14th afternoon shift.  I will suggest to the ACS referent person some improvements in the alarms criteria such that the DMS can recognize and alert on similar kind of plant failures.

Concerning the effect of a different working point of the temperature in the Laser Lab after the workaround, in Fig3 is reported a longer trend of IMC power, were it is possible that it  decreased as consequence of the ACS failure.

Images attached to this comment
39549_20170922173511_injfailure14sept2017imcpower2.gif 39549_20170922173517_injfailure14sept2017imcpower2zoom.gif 39549_20170922173524_imctrapowertrend.gif
derossi, genin - 18:22, Friday 22 September 2017 (39552)

The decreasing of the IMC transmitted power is probably a consequence of  the decreasing of the pre-mode cleaner transmitted power, due to a misalignment of the cavity linked to the temperature.

Images attached to this comment
39552_20170922181931_pmclosspower.png
Interferometer_sensitivity_studies (General)
Print this report.
bader - 10:18, Friday 22 September 2017 (39543)Get code to link to this report
Comment to Tapping tests in DET Lab (Click here to view original report: 39373)
Below are the noise budget plots for tapping tests on different mini links

1) North link (1188554325), large shoulder up to about 100 Hz
2) North link - 1 (1188555080
3) North link - 2 west (1188555153), structures excited at 300 - 400 Hz
4) North link - 2 east (1188555251, touches north link - 3
5) North link - 3 (118855404), huge effect, large shoulder up to 400 Hz

The NB plots lead to the conclusion that scattered light form the SDB2 bench is fed back to the detector when exciting the B1 mini link.
Images attached to this comment
39543_20170922101730_nlnew.png 39543_20170922101751_nl1.png 39543_20170922101815_nl2w.png 39543_20170922101824_nl2e.png 39543_20170922101832_nl3.png
AdV-DAQ (Data collection)
Print this report.
masserot - 08:51, Friday 22 September 2017 (39539)Get code to link to this report
DAQ troubles

Since the 20170920 there is(first plot):

  • some jumps on the latency  on at FbmFE level
  • and some holes in the data streams

At 20170921-13h55UTC I 've restarted all the Fbs servers, this fixed the holes problems only.

Looking more deeply on the latency Daq channels it apperars that the latency jumps was triggered by the SMS collection of the EarlyWarning servers done by the FbsMoni server(2nd plots).

I stopped and restarted  by hand the EarlyWarning server at 20170927-19h39UTC without any improvements. According the logfiles the EarlyWarning server was restarted again at 20h55UTC and after that there is no more jumps in the FbsMoni latency (3rd plot)

Images attached to this report
39539_20170922083940_20170922daqtrendlatencylevel1.png 39539_20170922083952_daqtroubles20170920.png 39539_20170922084556_daqtroubles20170920zoom.png
AdV-DET (Commissioning)
Print this report.
bersanetti, bonnand, casanueva, gouaty, kasprzack, valera - 22:31, Thursday 21 September 2017 (39537)Get code to link to this report
Investigations on OMC angular noise coupling and best SDB1 angular set points

Below is the list of actions performed during the shift of yesterday afternoon (Sep 20). All measurements were taken in low noise 3.

1/ Angular noise injections on SDB1, with bench at "nominal" position:

A first set of low frequency noise injections though the SDB1 actuators was performed while the bench was at its so-called "nominal" position, that is to say the angular position given by the B5 beam drift control, which roughly corresponded to the set point TX = -410 urad and TY = -3559 urad. The list of injections is reported in the table below.

Bench DoF Cut-off Amplitude (in Correction) Start (utc) Duration (s)
SDB1 TX 40 Hz 0.02 V RMS 14:49:33 220
SDB1 TX 40 Hz 0.08 V RMS 14:53:14 480
SDB1 TX 12 Hz 0.2 V RMS 15:02:34 360
SDB1 TY 12 Hz 0.2 V RMS 15:16:18 60
SDB1 TY 12 Hz 0.4 V RMS 15:17:29 60
SDB1 TY 12 Hz 0.8 V RMS 15:18:54 420
SDB1 TZ 12 Hz 0.2 V RMS 15:29:44 80
SDB1 TZ 12 Hz 2 V RMS 15:31:09 300

2/ Scan of the SDB1 set points (TX, TY)

Then we scanned the SDB1 bench set points by +/-5 um around the initial values (obtained with B5 drift control engaged).
 
2.1/ Scan in TX:
 
SDB1_LC_TX  = -410 ==> 15:50:00 UTC  (B5 drift control disabled in TX)
                            = -411 ==> 15:59:30 UTC
                            = -412 ==> 16:00:36 UTC
                            = -413 ==> 16:01:47 UTC
                            = -414 ==> 16:02:26 UTC
                            = -415 ==> 16:03:03 UTC
                            = -405 ==> 16:10:02 UTC
                            = -410 ==> 16:19:22 UTC
                            = -412 ==> 16:25:54 UTC
 
2.2/ Scan in TY:
 
During the scan in TY, the TX setpoint was kept at the value -412 urad which seemed to maximize the optical gain (see explanation below). For this reason the B5 drift control loop had to be kept disabled in TX.
 
SDB1_LC_TY  = -3559 ==> 16:28:10 UTC  (B5 drift control disabled both in TY and TX)
                            = -3554 ==> 16:32:24 UTC
                            = -3564 ==> 16:37:09 UTC
                            = -3558 ==> 16:47:20 UTC
                            = -3557 ==> 16:51:12 UTC
                            = -3558 ==> 16:52:50 UTC
                            = -3557.5 ==> 16:53:47 UTC
                            = -3558 ==> 16:55:29 UTC
                            = -3558.5 ==> 16:58:38 UTC
 
Please be aware that there is a transition phase between setpoints that last between 100s and 10s. Here the reported times are the ones when the command to change the setpoint to the requested value is launched (so this when the ramp starts towards the requested set point).  It might be wise to have a look a the data SDB1_LC_TY/TX when analysing these times.
 
During both scans in TX and TY we checked how the height of the NE and WE calibration lines seen in DARM (in particular the lines at 358.3 Hz and 355.8 Hz) were changing, which can give an indication of the interferometer optical gain. During the scan in TX, we noticed  a small improvement of the ITF response (height of the calibration lines higher by ~2%) for the position TX = -412 urad compared to the initial position (TX = -410 urad). Therefore we decided to keep this position as new working point. During the scan in TY, we also observed some slight variation of the calibration lines but here the optimal position was not so clear, and in any case closer to the initial value.
 
After these observations we added in vpm the possibility to add an offset on the B5 quadrants used to make SDB1 follow the B5 beam :
- SDB1_B5_QD2_H/V_pre       are the raw data from the quadrants.
- SDB1_B5_QD2_H/V_offset   are the offset one can set to change the bench alignment wrt to B5 (in SDB1_Quadrants config file).
- SDB1_B5_QD2_H/V               are the error signals used by the B5 drift control loop, the loop makes it to be 0.

Before the changes SDB1_B5_QD2_H/V were the raw data and error signals of the B5 drift loop as the offset channel did not exist.
Then after relock at LOW_NOISE_3 around 18:20:00 UTC, the vertical offset was set to 25 um (meaning the beam was shifted by -25 um on B5_QD2 vertically, SDB1_B5_QD2_V_pre = -25 um, which corresponded to the bench set point TX = -412 urad). Except for a few minutes between 18:39:25 UTC and 18:43:00 UTC during which the offset was set to 0 um (so the nominal condition for the alignment of SDB1).
 
3/ Angular noise injections on SDB1, with 25 um offset on B5_QD2_V
 
We performed two types of broad band noise injections with the SDB1 bench set at the new set point (with the 25 um offset in B5_QD2_V): some injections at high frequency (flat noise above 100 Hz) and some injections at low frequency (below 12 Hz).
 
3.1/ Noise injections above 100 Hz
 
Bench DoF Amplitude (in Correction) Start (utc) Duration (s)
SDB1 TX 5.5e-5 V RMS 18h58m31 120
SDB1 TX 2.7e-4 V RMS 19h00m40 100
SDB1 TX 2.7e-3 V RMS 19h02m23 180
SDB1 TX 1e-2 V RMS 19h05m48 300
SDB1 TX 5e-2 V RMS 19h11m03 140
SDB1 TY 5.5e-3 V RMS 19h23m52 60
SDB1 TY 2.2e-2 V RMS 19h24m55 120
SDB1 TY 5.5e-2 V RMS 19h27m14 210
 
3.2/ Noise injections below 12 Hz
 
Bench DoF Amplitude (in Correction) Start (utc) Duration (s)
SDB1 TX 0.2 V RMS 19:37:35 160
SDB1 TX 0.1 V RMS 19:40:18

180

 
 
4/ Angular noise injections on SDB1 with large misalignment
 
At 20:00:00 utc we changed the offset in B5_QD2_V from -25 to -180 um which moved the bench from the position TX = -413 to TX = -419. This made the ITF lock quite unstable and Diego had to adjust offsets in the alignment loops to compensate this SDB1 misalignment. We then started to inject a low frequency noise (cut-off at 12 Hz) in TX at 20:26:49 utc, amplitude 0.1 V RMS. The noise injection lasted a few minutes until the ITF unlocked.
 
We stopped the activity here. We decided to put back the offset of 25 um in the B5_QD2_V signal of SDB1 and leave it like this. It is not clear if this offset (which was found to correspond to the optimal position of SDB1 in TX for what concerns optical gain maximization) is going to be stable in time. Therefore, one useful test to do in one of the next shifts would be to perform a scan of the OMC1 temperature with the beam directly reflected from north arm in order to check that the OMC alignment is still correct with this offset.
 
Analysis of these data to be followed.
Comments related to this report
bersanetti, ruggi, gouaty, bonnand, frolov, kasprzack - 19:43, Friday 22 September 2017 (39554)
We wanted to investigate if the 270 Hz and nearby noise peaks are from the OMC beam jitter.
The first plot compares the best reference from last week (when the 270 peaks were very low) to the current typical noise floor and to the two bench misalignments (TX, TY). In one of the misaligned state (TY), the 270 peak height increased as if it was driven by the beam jitter. But we didn't observe the same increase for the other misalignments - TX or TY.
The second plot shows the response to excitations from 100 to 400 Hz. A different set of peaks was excited while the height of 270 peaks remains unaffected. This would mean that we are not only looking at OMC beam jitter coupling, but also at some other noise coupling that gets affected by the global alignment (by moving SDB1, we are also moving the beam on SDB2, which has QPDs used for common arm alignment).

Furthermore, the DARM coherence with PRCL at 270 Hz is high according to BRUCO (see third plot). This makes us suspect that the peak is actually from frequency noise coupling, and it gets modulated by the global alignment.

The 270 peak was persistently in the DARM spectrum starting from September 15. Prior to this date, there were several days when the best spectrum of the day didn't have the peak. Coincidentally, the injection lab AC status changed on September 14.

During the same shift, we tried to see if SPRB alignment had an effect on the peaks, but scanning the full SPRB_B4_112 and LSC_B4_DC ranges didn't modulated the 270 peak.
Images attached to this comment
39554_20170922193531_sdb1tyoffset.png 39554_20170922193539_sdb1exc.png 39554_20170922193548_lscprcl.png
genin, fiori - 11:01, Saturday 23 September 2017 (39564)

Regarding the 270 Hz peak which seems to be present on DARM since the Laser room to SIB2 room door and cleanroom to SIB2 door have been opened. We can clearly see it on the 

the Beam Pointing control system signals (see attached plot). We could not see any excess of accoustic noise at that frequency but there is a slight effect on the seismic noise measured with the accelerometer on top of the EIB which suggest that a resonance of a mount (maybe EIB_M6 one) could be excited (see entry https://logbook.virgo-gw.eu/virgo/?r=38640).

The strange thing is that the excess of noise read by the BPC quadrants is very small and we are not expecting any limitation from input beam jitter noise at that level when looking at the last noise projection we made on the 13th of september (https://logbook.virgo-gw.eu/virgo/?r=39459). At that time, we injected a large boradband noise without being able to see any coherence for frequencies higher than 100Hz.

Maybe this noise comes from the CB clean rooms and couple the ITF through PRCL (SIB2 PDs)?

We plan to make more investigations next week on that issue trying to understand it better and solve it.

Images attached to this comment
39564_20170923110048_plot3.gif
AdV-COM (AdV commissioning (1st part) )
Print this report.
fiori, pasqualetti, chiummo, vacuum team - 17:59, Thursday 21 September 2017 (39538)Get code to link to this report
Stiffening of SPRB mini link

This morning we added a mechanical structure to the pipe linking the PR tower to the SPRB minitower with the aim to make it more stiff and possibly increase the frequency of its first modes.
The structure consists of a 'strong' leg bolted to the PR tower base and to the link pipe itself, and it mainly acts in the vertical direction (see attachements 1,2 and 3).
It can be removed when necessary. Note that the sand bag added on Sept 5th was left unchanged.

We then performed a tapping of the pipe structure (ITF unlocked) to check the mechanical modes.
Tapping along the radial directions (X and Y) excite mostly frequencies above 300Hz, while tapping along the beam direction (Z) excites mainly one mode around 200Hz  (see attachements 4, 5 and 6).
Overall the machanical response of the pipe seems improved with respect to before: the tests done on May 23 showed radial mechanical modes of the pipe around 55Hz  and 60Hz which also showed up in Z.
Presently the first modes look shifted up in frequencies and better dumped. The quiet spectra (last attachement) show a peak at 60Hz that moved to about 70Hz with reduced amplitude (X) and a peak at about 55Hz that moved to about 80Hz also with reduced amplitude (Y and Z).
Note that this comparison also includes the sand bag, more analysis to be done to disentangle the effect. Effects on Hrec and B4 beam to be checked.

Attached:

  1. picture
  2. picture
  3. tapping along X (radial horizontal direction: tapping on the pipe between the leg and PR tower)
  4. tapping along Y (radial vertical direction: tapping location as above)
  5. tapping along Z (beam direction: tapping acting on the flange)
  6. quiet spectra compared before and after the leg (plus the sand bag)
  7. drawings
Images attached to this report
39538_20170921175829_img20170921113753.jpg 39538_20170921175833_img20170921113826.jpg 39538_20170921175848_tapx.gif 39538_20170921175856_tapy.gif 39538_20170921175902_tapz.gif 39538_20170921175908_primadopoxyz.gif
Non-image files attached to this report
Detector Operation (Operations Report)
Print this report.
menzione - 16:00, Thursday 21 September 2017 (39535)Get code to link to this report
Operator Report - Morning Shift 21/09

Upon arrival I found ITF Locked in Low_Noise_3 with Horizon ~30 Mpc.

At 06:30 UTC I set ITF in Commissioning Mode.
From 06:30 to 08:46 UTC the planned Magnetic Injections in MC Building, carried out by Fiori and Cirone, went on without problems.
Upon concluded ITF unlocked and the following activities were performed:

  • Planned SPRB stiffening (carried out by Vacuum crew).
  • Cleaning of the new NI marionette in Cleaning Lab (carried out by M. Ciardelli, F. Gherardini and Roma1 team).
  • Infrastructure checks and measurements (carried out by A. Paoli and A. Buggiani).
  • SR / SPRB hardware installation (carried out by V. Dattilo, R. Passaquieti and V. Boschi). For this activity Boschi opened SPRB and SR Local Controls.
  • Pillant went in atrium (LL) for a check.

All these morning activities are concluded at 09:45 UTC except for the SR / SPRB hardware installation that required the replacement of a board on SR Rack. It has been replaced and restarted only at 12:45 UTC.

ITF relocked at second attempt at 13:00 UTC with Horizon ~29 Mpc.
The planned activity on tapping test on SPRB with three new accelerometers has been postponed for tomorrow morning. People involved I. Fiori and A. Chiummo.

DMS - Infrastructure
At 06:20 UTC, under request of D. Soldani, I changed the threshold of the INF_TB_CW_TANK_TE (Infrastructures -> ACS_TB) from 5 to 4 degrees. InfraMoniAlarmed properly restarted from VPM.

AdV-DAQ (Data Acquisition and Global Control)
Print this report.
masserot - 15:02, Thursday 21 September 2017 (39534)Get code to link to this report
Comment to Monitoring the NTP time and the GPS (Click here to view original report: 39515)

at  Sep 21 14:37:31 LT the rtpc5 ntp daemon was restarted to use ntp server ntp2.inrim.it instead of : ntp.obspm.fr previously , the goal is to have 2 rtpcs using the same ntp server:

  • channel LNFS_NTP_delta_gps on rtpc6 with the ntp server: ntp2.inrim.it, ACL server: LNFS_ntp
  • channel SDB2_NTP_delta_gps on rtpc5 with the ntp server: ntp2.inrim.it, ACL server: SDB2_ntp
AdV-COM (AdV commissioning (1st part) )
Print this report.
cirone, fiori - 11:34, Thursday 21 September 2017 (39528)Get code to link to this report
Magnetic injections at MCB
Yesterday afternoon we prepared the setup for magnetic injections inside mode cleaner building, that we performed this morning between 08:30 and 10:30. We injected magnetic noise lines between 14 Hz and 140 Hz by sending 0.5 V (1V pkpk) signals lasting 300 sec each to the amplifier connected with the coil (axes orientation W-E) with Onosokky generator. The magnetic activity is monitored with the uniaxial magnetometer already present in the building (N-S: channel to be corrected because currently is set to W-E) and one tri-axial magnetometer (see photos for the locations).
Apparently the injected lines are not visible in Hrec and so We will probably get only an upper limit for the magnetic projection.

Log file:
quiet 1190013616 1190013686

Freq gpsstart gpsend
14 1190013758 1190014282
26.6 1190014324 1190014630
39.2 1190014672 1190014980
51.8 1190015003 1190015310
64.4 1190015340 1190015650
77 1190015670 1190015980
89.6 1190016009 1190016320
102.2 1190016360 1190016670
114.8 1190016720 1190017030
127.4 1190017080 1190017400
140 1190017450 1190017760


Images attached to this report
39528_20170921113322_img20170921084020952.jpg 39528_20170921113335_img20170920143243506.jpg 39528_20170921113342_img20170920143252734.jpg 39528_20170921113357_sensorcoilposition.png 39528_20170921113405_img20170921094039979.jpg
Detector Operation (Operations Report)
Print this report.
Montanari - 22:57, Wednesday 20 September 2017 (39527)Get code to link to this report
Operator Report - Afternoon Shift 20/09

All the afternoon shift was dedicated to DET activity, wich finished at 20:50UTC.

At 17:00UTC an earthquake in Japan, magnitude 6.2, unlocked ITF and prevented the locking until 18:00UTC. Finally, at 18:22UTC ITF relocked in LOW_NOISE_3.

20:50:UTC - Stopped Commissioning Mode.
ITF left locked in LOW_NOISE_3

DAQ
18:16UTC, 18:19UTC, 19:09UTC missing data

AdV-COM (AdV commissioning (1st part) )
Print this report.
cirone, fiori - 16:53, Wednesday 20 September 2017 (39525)Get code to link to this report
New Magnetic noise injections at CEB
Yesterday morning (19/09/17) We moved the coil from the highest terrace to the east terrace at level 4 (see photos), in order to repeat the injections done in July (38771; 38369). In addiction to the Three uniaxial magnetometers in L-room, we added two more triaxial probes, near NI and Wi towers (sampled at 1kHz), in oder to make the projection in three different locations inside the building (photo 7). The coil axes is oriented along N-S direction.
During the night we performed the injection, starting from roughly 10:35 pm, dividing it in two cycles of 12600 seconds each (7 hours in total). The amplitude of the signal sent to the coil was 0.6 V. We injected 21 equally spaced lines from 14 Hz to 140 Hz.

Logfile:

% NOISE_CEB_DAQroom
% Cycle 1 of 2. Measure time: 12600.0 seconds
% gpsstart duration(s) gpsstop frequency(Hz) amplitude(V)
1189888475.55 600.0 1189889075.64 14.0 0.6
1189889075.64 600.0 1189889675.71 20.3 0.6
1189889675.72 600.0 1189890275.82 26.6 0.6
1189890275.82 600.0 1189890875.91 32.9 0.6
1189890875.91 600.0 1189891476.0 39.2 0.6
1189891476.0 600.0 1189892076.08 45.5 0.6
1189892076.08 600.0 1189892676.18 51.8 0.6
1189892676.18 600.0 1189893276.25 58.1 0.6
1189893276.25 600.0 1189893876.29 64.4 0.6
1189893876.29 600.0 1189894476.39 70.7 0.6
1189894476.39 600.0 1189895076.49 77.0 0.6
1189895076.49 600.0 1189895676.55 83.3 0.6
1189895676.55 600.0 1189896276.65 89.6 0.6
1189896276.66 600.0 1189896876.76 95.9 0.6
1189896876.76 600.0 1189897476.86 102.2 0.6
1189897476.86 600.0 1189898076.94 108.5 0.6
1189898076.94 600.0 1189898677.04 114.8 0.6
1189898677.04 600.0 1189899277.14 121.1 0.6
1189899277.14 600.0 1189899877.24 127.4 0.6
1189899877.24 600.0 1189900477.34 133.7 0.6
1189900477.34 600.0 1189901077.38 140.0 0.6
% NOISE_CEB_DAQroom
% Cycle 2 of 2. Measure time: 12600.0 seconds
% gpsstart duration(s) gpsstop frequency(Hz) amplitude(V)
1189901077.39 600.0 1189901677.49 14.0 0.6
1189901677.49 600.0 1189902277.57 20.3 0.6
1189902277.57 600.0 1189902877.66 26.6 0.6
1189902877.66 600.0 1189903477.76 32.9 0.6
1189903477.77 600.0 1189904077.87 39.2 0.6
1189904077.87 600.0 1189904677.97 45.5 0.6
1189904677.97 600.0 1189905278.07 51.8 0.6
1189905278.07 600.0 1189905878.17 58.1 0.6
1189905878.17 600.0 1189906478.27 64.4 0.6
1189906478.27 600.0 1189907078.34 70.7 0.6
1189907078.34 600.0 1189907678.44 77.0 0.6
1189907678.44 600.0 1189908278.52 83.3 0.6
1189908278.52 600.0 1189908878.62 89.6 0.6
1189908878.63 600.0 1189909478.72 95.9 0.6
1189909478.73 600.0 1189910078.83 102.2 0.6
1189910078.83 600.0 1189910678.93 108.5 0.6
1189910678.93 600.0 1189911279.03 114.8 0.6
1189911279.03 600.0 1189911879.13 121.1 0.6
1189911879.13 600.0 1189912479.22 127.4 0.6
1189912479.23 600.0 1189913079.33 133.7 0.6
1189913079.33 600.0 1189913679.42 140.0 0.6
Images attached to this report
39525_20170920162313_img20170920105248036.jpg 39525_20170920162333_img20170920105258441.jpg 39525_20170920162459_img20170919124309004.jpg 39525_20170920162521_img20170919124324842.jpg 39525_20170920162541_img20170919121750879.jpg 39525_20170920162616_img20170919121728672.jpg 39525_20170920164538_posizionesensoricoil.png
AdV-COM (AdV commissioning (1st part) )
Print this report.
ruggi, bersanetti, Frolov, Kasprzack - 16:50, Wednesday 20 September 2017 (39526)Get code to link to this report
change of PR_F7_TY (POP) set point

After the adjustment of POP set point performed on Sep 08, the position of the beam on B4 camera has never checked anymore. Today we found the beam out of the standard position on the camera. We put it back moving the POP about in the position it was before the adjustment of Sep 08 (Sc_PR_F7_TY = - 112).

AdV-COM (AdV commissioning (1st part) )
Print this report.
ruggi - 16:21, Wednesday 20 September 2017 (39524)Get code to link to this report
Comment to Many changes in suspension and alignment control strategy (Click here to view original report: 39493)

An undesired side effect of GIPC has been the disabling of fModErr loop (the tidal control of common arm length which keeps MC length tuned on the modulation frequency). This means that since a few days we are working with fModErr slightly mistuned. It cannot be excluded some impact of this defect on the sensitivity. The bug will be fixed as soon as possible.

AdV-SAT (Suspension control upgrade)
Print this report.
Boschi - 16:13, Wednesday 20 September 2017 (39523)Get code to link to this report
MC IP Recentered
MC IP has been recentered using motors in order to decrease the top stage actuation DC values since they were very close to saturation as shown in figure.
Images attached to this report
39523_20170920161350_mccoils.png